# New Approach to Strong Correlation: Twisting Hubbard into the Orbital Hatsugai-Kohmoto Model

### 16 Jan 2024 v1

@@ Abstract
Based on the surprising observation that all the correlation functions in the local-in-momentum space Hatsugai-Kohmoto (HK) model fall off exponentially with distance, we show that the $N_{\alpha}$-orbital extension has a fundamental connection with the Hubbard model. Namely, for $N_{\alpha}$-orbitals on each of the $N$ ${\bf k}$-points and dynamics between the orbital degrees of freedom, the $N_{\alpha}$-orbital HK model is equivalent to covering the Brillouin zone with $N/N_{\alpha}$ Hubbard clusters each containing $N_{\alpha}$-sites all connected via twisted boundary conditions. The thermodynamic limit arises from the ${\bf k}$-state starting point of the HK model. While the Hubbard model necessarily results when $N_{\alpha}=N$, we show that already at $N_{\alpha}=2,4$, standard results for the Hubbard model emerge, such as the onset of an insulating state regardless of the strength of the interactions, double occupancy, dynamical spectral weight transfer, charge neutral excitations leading to the algebraic temperature dependence of the specific heat, all with a fraction of the computational time of more advanced cluster methods and making it possible to obtain analytical insights. Consequently, the $N_{\alpha}$-orbital HK model offers a new tool for strongly correlated quantum matter.
@@ II Local Correlations
The starting point for our analysis is Fig. 1a which shows that although the HK model has all-to-all interactions, the 2-point correlator falls off exponentially. The same is true for the 4-point and all higher correlation functions. This is significant because Wightman’s reconstruction theoremWightman (1956) ensures that if all the n-point functions are local, then so is the theory. The question arises, why does a model admit non-local interactions at the level of the Hamiltonian but exhibit purely local correlations? To gain insight into this, we plot in Fig. 1b the density of states for various $U$ across the Mott transition of the half-filled band HK model, which interestingly has never been computed. Here the Mott transition to the insulating state proceeds via depletion of the density of states at the central peak at $\omega=0$ as in the standard single-site DMFT modelGeorges et al. (1996). This is telling and consistent with the exponential fall-off of the n-point correlators: all the physics is local regardless of the non-local nature of the Hamiltonian. This can be understood simply. The Hamiltonian for the band HK model is identical in form to the atomic limitManning-Coe and Bradlyn (2023) of the Hubbard model with the band index ${\bf k}$ replacing the site index ${\bf i}$. A more precise statement will be made later when we complete the mapping to the full Hubbard model.
Figure 2: Scheme of the orbital unit cell (encircled by red lines) and lattice (dashed lines) for the orbital HK model: a) 1-orbital= band HK model. b) The 2-orbital HK model leads to a doubling of the unit cell size and consequently halves the Brillouin zone (BZ). c) The 4-orbital HK model results in a quadrupling of the unit cell size and reduces the BZ to one-quarter of its original size.
Figure 3: Density of states representing Mott transition at half-filling of the (a) 2-orbital and (b) 4-orbital HK model (sharing the same legend), and (c) double occupancy at half-filling ($x=0$) as a function of $U$ for the orbital HK model with various $N_{\alpha}$. The green dots are from auxiliary-field quantum Monte Carlo (QMC) simulations on the 2D Hubbard modelXu et al. (2024). The inset of panel (c) compares the 4-orbital HK and QMC resultsXu et al. (2024) at $1/8$-hole-doping ($x=0.125$).
@@ III Orbital HK
However, just like DMFT, the basic band HK model does not capture the full Hubbard physics. As pointed out previouslyManning-Coe and Bradlyn (2023), band HK can be generalized to include multiple orbitals per ${\bf k}$-state as is necessary for the adaption of this model to QAH and QSH topological models, all of which have 2 atoms per unit cellMai et al. (2023b, c). Such a change produces qualitatively new physics as the hybridization between the orbitals liftsManning-Coe and Bradlyn (2023) the thermodynamic degeneracy of the band HK model. Inspired by that, we adopt an orbital HK model to capture the strongly correlated dynamical mixing in a simple square lattice by redefining a set of lattice sites as orbitals. The most general version of this model is given by
$H_{\text{OHK}}=&\sum_{{\bf k},\alpha,\alpha^{\prime},\sigma}g_{\alpha,\alpha^{\prime}}({\bf k})c_{{\bf k}\alpha\sigma}^{\dagger}c_{{\bf k}\alpha^{\prime}\sigma}-\mu\sum_{{\bf k},\alpha}n_{{\bf k}\alpha,\sigma}\\ &+\sum_{{\bf k},\alpha,\alpha^{\prime}}U_{\alpha,\alpha^{\prime}}n_{{\bf k}\alpha\uparrow}n_{{\bf k}\alpha^{\prime}\downarrow},\end{aligned}$ (1)
where $g(\bf k)$ is the dispersion matrix determined by the underlying lattice hopping process and orbital setup, $\alpha(\alpha^{\prime})$ is the orbital index and $\sigma$ is the spin. The computational time is governed by the diagonalization of the Fock-space matrix of dimension $4^{N_{\alpha}}\times 4^{N_{\alpha}}$. The redefinition of lattice sites as $N_{\alpha}$ orbitals is depicted in Fig. 2. Explicitly shown are the 2 and 4-orbital cases in Fig. 2(b) and (c) respectively. Note that the orbital interactions are all local in the orbital basis while the hoppings occur among different orbitals, hence representing the built-in Hubbard physics for an $N_{\alpha}$-site clusterManning-Coe and Bradlyn (2023). Also, we have included inter-orbital interactions in the most general case while our simulations shown below are only for the intra-orbital case. To see the effect of just a 2-orbital extension, look no further than the corresponding density of states at half-filling in Fig. 3a. Any non-zero $U$ is sufficient to eliminate the metallic state producing a vanishing of the density of states at zero energy. Also of note is the narrowing of the bands which is completely absent in the band HK model. No qualitatively new physics appears in the Mott transition of the 4-orbital HK model at half-filling (Fig. 3b). These results are similar to those from state-of-the-art cluster versions of DMFT methodsMaier et al. (2005); Georges et al. (1996). We also calculate the double occupancy defined as
@@
For the band HK model, half the Brillouin zone (BZ) is doubly occupied. That is, $\langle n_{k\alpha\sigma}\rangle=1$ if occupied and zero otherwise. Consequently, $D_{\rm non-int}=1/2$. On the other hand, for the orbital model, $\langle n_{k\alpha\sigma}\rangle=\langle n_{k\beta\sigma}\rangle$ for all $\alpha\neq\beta$ by rotational symmetry. In addition, $\langle n_{k\alpha\sigma}\rangle=1/2$ in the reduced Brillouin zone is not unity as in the starting band HK model. As a result, we find that for the orbital HK model, $D_{\rm non-int}=1/4$. Note that this result is identical to the non-interacting limit of the Hubbard model and underscores how drastically band HK differs from its orbital counterpart. The plots for the $D$ show this dramatic difference. In band HK, $D$ decreases steadily from $1/2$ and vanishes for $U>W$. However, in orbital HK model, $D$, starting at 1/4, just tapers asymptotically as $U$ increases. Note also the rapid convergence between the 2 and 4-orbital cases which matches perfectly with state-of-the-art auxiliary-field quantum Monte Carlo simulaltions (QMC) on the Hubbard modelXu et al. (2024). The inset also shows the double occupancy of the 4-orbital HK model at $1/8$-hole-doping ($x=0.125$) and its comparison with QMC simulationsXu et al. (2024). Once again, the agreement is remarkable underscoring that orbital HK converges rapidly to Hubbard physics. Additionally, the computational cost for solving the 4-orbital HK model is considerably lower, both in terms of algorithm development (40 lines of Mathematica) and actual code execution, when compared to QMC or cluster-type DMFT simulations on the Hubbard model. Most crucially, the orbital HK model is more amenable to analytical calculations, an aspect we intend to explore in future studies.
The fact that the double occupancy only vanishes asymptotically as $U$ increases in the orbital HK model implies the presence of dynamical mixing between the upper and lower Mott sub-bands. This arises entirely from the non-commutativityHarris and Lange (1967); Meinders et al. (1993); Eskes et al. (1991) of the kinetic and potential energy and appears as $t/U$ corrections to the low-energy spectral weight (LESW). In the atomic limit, the LESW is strictly $2x$ because each hole can be occupied by a spin-up or spin-down electron. Any dynamical corrections to this necessarily generate double occupancy and hence increase the LESW strictly defined as
@@
In Fig. (4), we compare the LESW of the 4-orbital HK model with Fig. (3) of Ref. [Meinders et al., 1993]. There is no qualitative difference with exact diagonalization on the (1D) Hubbard modelMeinders et al. (1993); Eskes et al. (1991). Both increase faster than $2x$. The semiconductor (dashed line) and Fermi liquid (dashed-dotted) results are shown for comparison. As the occupied part of the lower band has a weight $1-x$, the total weight in the lower band now exceeds $1+x$. As only $1+x$ electrons can occupy the lower band, dynamical spectral weight transfer (DSWT) defined as $\Lambda(x)-2x$ and plotted in Fig. (4b) impliesPhillips (2010) that the spectral weight in the lower band cannot be exhausted by counting electrons alone. Fig. (4b) is quantitatively in agreement with the Hubbard model (Fig. (4) of Ref. [Meinders et al., 1993]), the only difference being the maximum value which for 4-orbital HK is $0.15$ whereas it is $0.23$ for the (1D) Hubbard result. Consequently, that the LESW in the orbital HK model increases faster than $2x$ is a profoundly non-trivial result as no analytically solvable model has ever been formulated to capture this feature of the Hubbard model.
Figure 4: LESW and DSWT in the exact solution of the 4-orbital HK model ($U=20$) for the hopping parameters shown in panels (a) and (b) respectively. The solid line shown with slope $2x$ is the band HK or atomic Hubbard result. The dashed and dotted lines depict the semiconductor and Fermi liquid results respectively. Note, there is no qualitative difference with the exact diagonalization results for the Hubbard modelMeinders et al. (1993); Eskes et al. (1991). As $t/U$ increases, so does the DSWT.
All of these results point to a fundamental connection with the Hubbard model. Fig. 2 contains the essential idea. Grouping the $N$ lattice sites (or ${\bf k}$-points) into cells (Fig. 2(a-c)) with $N_{\alpha}$ orbitals reduces the number of unit cells (or ${\bf k}$-points) to $N/N_{\alpha}$. That is, the high momenta are systematically being removed. Each cell in momentum space now contains the exact information of the $N_{\alpha}$-site Hubbard model. For example, $N_{\alpha}=2$ signifies that the orbital basis contains the information of the Hubbard dimer which has a non-degenerate ground state, and hence the spurious $2^{N}$ spin degeneracy of the band-HK model is removed. In addition to hybridization between orbitals in each cell, the cells are also connected in terms of hopping and interactions. For $N_{\alpha}=4$, we show the explicit connection between the cells. The matrix $g({\bf k})$, representing the hopping connectivities in Fig. 2c, is given by
@@
for the 4-orbital case, with $\varepsilon_{tx}=-2t\cos k_{x}$, $\varepsilon_{ty}=-2t\cos k_{y}$ and $\varepsilon_{t^{\prime}}=-4t^{\prime}\cos k_{x}\cos k_{y}$, identical in form to that in the $2\times 2$ Hubbard cluster with twisted boundary conditions (TBC) Huang (2020) where $\alpha$ is replaced with the site index $i$ and $k_{x}=\theta_{x}$ and $k_{y}=\theta_{y}$ where $\theta_{x}$ and $\theta_{y}$ define the TBC. Besides, TBC gives local-in-momentum interactions in the reduced BZ which is exactly the interaction term in Eq. (1). For any $N_{\alpha}<N$, the orbital HK model corresponds to a $N/N_{\alpha}$ copies of $N_{\alpha}$-site Hubbard clusters all connected by TBC. The matrix that determines the TBC is purely local as it is determined by the dynamics. Hence, orbital HK is purely local in the reduced Brillouin zone regardless of the number of orbitals used, even $N_{\alpha}=1$ corresponding to band HK whose locality is illustrated in Fig. 1. In this case, the TBC matrix is simply $\varepsilon_{\bf k}$. This exact mapping is a surprise and offers a new way of interpreting exact diagonalization with TBC. The most general statement is as follows. Theorem: The orbital HK model is equivalent to covering the Brillouin zone with $N/N_{\alpha}$ Hubbard clusters each containing $N_{\alpha}$-sites all connected via TBC. Clearly, the $N$-site Hubbard model results when $N_{\alpha}=N$. Loosely, the $N/N_{\alpha}$ identical copies throughout the Brillouin zone can be thought of as replica copies.
@@
Aside from DSWT, the orbital HK model also exhibits a pseudogap in the density of states. Shown in Fig. 5 is the density of states for the 4-orbital HK model with (Fig. 5(a,c)) and without (Fig. 5(b,d)) the next-nearest-neighbor hopping, $t^{\prime}$. Fig. 5(a) shows a suppression of the density of states at the chemical potential ($\rho(\omega=0)\lesssim 0.1$) in the underdoped region with $t^{\prime}=-0.25$ when $U\geq W$ ($W=8t$ is the bare bandwidth), indicative of a pseudogap in the absence of superconducting order. The complete tracking of the density of states at zero frequency in Fig. 5(c) displays the trend that as $U$ increases, the pseudo-gap region appears and extends to higher hole-doped density. In contrast, when $t^{\prime}=0$, the pseudo-gap suppression does not arise until $U\gg W$, as shown in Fig. 5(b,d). These observations are consistent with a recent studyHuang (2020) on a $2\times 2$ Hubbard cluster with TBC. Finally, we compute the heat capacity of the 4-orbital HK model at half-filling shown in Fig. 6. Most noticeable is the two-peak structure at $U>W$ representing a demarcation of the charge and spin degrees of freedom into high and low-temperature regimes, respectively, consistent with QMC simulations on the Hubbard model Duffy and Moreo (1997); Wang et al. (2022). Also at half-filling, we find a near-crossing of the heat capacity curves as a function of $U$ at a temperature intermediate between the spin and charge excitations. The Maxwell relations governing the entropy dictateVollhardt (1997) such a crossing. Since there is no sign-problem restriction, we can access the low-temperature heat capacity exactly. As we show in Fig. 6(b), the low-temperature heat capacity data follow a power-law increase detailed in the supplement. While such algebraic growth might seem counterintuitive for a Mott insulator, the first-excited state is charge neutralMai et al. (2024); Zhao et al. (2023a) as shown in the supplement. Such charge-neutral excitations determine the low-T behavior of the specific heat.

### 23 Jan 2024 v2

@@ Abstract
Based on the surprising observation that all the correlation functions in the local-in-momentum space Hatsugai-Kohmoto (HK) model fall off exponentially with distance, we show that the $n$-orbital extension has a fundamental connection with the Hubbard model. Namely, the $n$-orbital HK model is equivalent to covering the Brillouin zone with $N/n$ Hubbard clusters each containing $n$-sites all connected via twisted boundary conditions. The thermodynamic limit arises from the ${\bf k}$-state starting point of the HK model. While the Hubbard model necessarily results when $n=N$, we show that already at $n=2,4$, standard results for the Hubbard model emerge, such as the onset of an insulating state regardless of the strength of the interactions, double occupancy, dynamical spectral weight transfer, charge neutral excitations leading to the algebraic temperature dependence of the specific heat, all with a fraction of the computational time of more advanced cluster methods and making it possible to obtain analytical insights. Consequently, the $n$-orbital HK model offers a new tool for strongly correlated quantum matter.
@@ II Local Correlations
The starting point for our analysis is Fig. 1a which shows that although the HK model has all-to-all interactions, the 2-point correlator falls off exponentially. The same is true for the 4-point and all higher correlation functions. This is significant because Wightman’s reconstruction theoremWightman (1956) ensures that if all the n-point functions are local, then so is the theory. The question arises, why does a model admit non-local interactions at the level of the Hamiltonian but exhibit purely local correlations? To gain insight into this, we plot in Fig. 1b the density of states for various $U$ across the Mott transition of the half-filled band HK model, which interestingly has never been computed. Here the Mott transition to the insulating state proceeds via depletion of the density of states at the central peak at $\omega=0$ as in the standard single-site DMFT modelGeorges et al. (1996). This is telling and consistent with the exponential fall-off of the n-point correlators: all the physics is local regardless of the non-local nature of the Hamiltonian. This can be understood simply. The Hamiltonian for the band HK model is identical in form to the atomic limitManning-Coe and Bradlyn (2023) of the Hubbard model with different site energies with the band index ${\bf k}$ replacing the site index ${\bf i}$. A more precise statement will be made later when we complete the mapping to the full Hubbard model.
Figure 2: a) Evolution of the reduced Brillouin zone (with the constraint $N\times n=N_{\rm sites}$) as the number of orbitals ($n$) increases, leading to a purely local (in k-space) model when $n=N_{\rm sites}$. Alternative real-space construction by grouping the atoms into cells (encircled by green/red lines) with an updated lattice constant. b) Original unit cell in the band HK model with primitive unit cell in dashed line. c) The 2-atom (orbital) HK model leads to a doubling of the unit cell size and the lattice constant (dashed line) d) The 4-orbital HK model results in a quadrupling of the unit cell size. When $n=N_{\rm sites}$, the unit cell contains all the sites.
Figure 3: Density of states representing Mott transition at half-filling of the (a) 2-orbital and (b) 4-orbital HK model (sharing the same legend), and (c) double occupancy at half-filling ($x=0$) as a function of $U$ for the orbital HK model with various $n$. The green dots are from auxiliary-field quantum Monte Carlo (QMC) simulations on the 2D Hubbard modelXu et al. (2024). The inset of panel (c) compares the 4-orbital HK and QMC resultsXu et al. (2024) at $1/8$-hole-doping ($x=0.125$).
@@ III Orbital HK
However, just like DMFT, the basic band HK model does not capture the full Hubbard physics. As pointed out previouslyManning-Coe and Bradlyn (2023), band HK can be generalized to include multiple orbitals per ${\bf k}$-state as is necessary for the adaption of this model to QAH and QSH topological models, all of which have 2 atoms per unit cellMai et al. (2023b, c). Such a change produces qualitatively new physics as the hybridization between the orbitals liftsManning-Coe and Bradlyn (2023) the thermodynamic degeneracy of the band HK model. Inspired by that, we adopt an orbital HK model to capture the strongly correlated dynamical mixing in a simple square lattice by redefining a set of lattice sites as orbitals. This will of course lead to a reduction in the size of the Brillouin zone (BZ), referred to as the reduced Brillouin zone (rBZ), as shown in Fig. 2a. The most general version of this model is given by
$H_{\text{OHK}}=&\sum_{k,\alpha,\alpha^{\prime},\sigma}g_{\alpha,\alpha^{\prime}}({\bf k})c_{{\bf k}\alpha\sigma}^{\dagger}c_{{\bf k}\alpha^{\prime}\sigma}-\mu\sum_{{\bf k},\alpha}n_{{\bf k}\alpha,\sigma}\\ &+\sum_{{\bf k},\alpha,\alpha^{\prime}}U_{\alpha,\alpha^{\prime}}n_{{\bf k}\alpha\uparrow}n_{{\bf k}\alpha^{\prime}\downarrow},\end{aligned}$ (1)
where $k$ is summed over the rBZ of the underlying lattice structure, $g(\bf k)$ is the dispersion matrix determined by the underlying lattice hopping process and orbital setup, $\alpha(\alpha^{\prime})$ is the orbital index and $\sigma$ is the spin. The redefinition of lattice sites as $n$ orbitals is depicted in Fig. 2(b-d). Explicitly shown are the 2 and 4-orbital cases in Fig. 2(c) and (d) respectively. While the interaction term as written includes both inter and intra-orbital repulsions, only the latter need be retained to recover the n-site Hubbard clusterManning-Coe and Bradlyn (2023). Our calculations below are only for the intra-orbital case. The computational time is governed by the diagonalization of the Fock-space matrix of dimension $4^{n}\times 4^{n}$. To see the effect of just a 2-orbital extension, look no further than the corresponding density of states at half-filling in Fig. 3a. Any non-zero $U$ is sufficient to eliminate the metallic state producing a vanishing of the density of states at zero energy. Also of note is the narrowing of the bands which is completely absent in the band HK model. No qualitatively new physics appears in the Mott transition of the 4-orbital HK model at half-filling (Fig. 3b). These results are similar to those from state-of-the-art cluster versions of DMFT methodsMaier et al. (2005); Georges et al. (1996). We also calculate the double occupancy defined as
@@
For the band HK model, half the BZ is doubly occupied. That is, $\langle n_{{\bf k}\alpha\sigma}\rangle=1$ if occupied and zero otherwise. Consequently, $D_{\rm non-int}=1/2$. On the other hand, for the orbital model, $\langle n_{{\bf k}\alpha\sigma}\rangle=\langle n_{{\bf k}\beta\sigma}\rangle$ for all $\alpha\neq\beta$ by rotational symmetry. In addition, $\langle n_{{\bf k}\alpha\sigma}\rangle=1/2$ in the reduced Brillouin zone is not unity as in the starting band HK model. As a result, we find that for the orbital HK model, $D_{\rm non-int}=1/4$. Note that this result is identical to the non-interacting limit of the Hubbard model and underscores how drastically band HK differs from its orbital counterpart. Fig. 3c for the $D$ shows this dramatic difference. In band HK, $D$ decreases steadily from $1/2$ and vanishes for $U>W$. However, in orbital HK model, $D$, starting at $1/4$, just tapers asymptotically as $U$ increases. Note also the rapid convergence between the 2 and 4-orbital cases which matches perfectly with state-of-the-art auxiliary-field quantum Monte Carlo simulations (QMC) on the Hubbard modelXu et al. (2024). The inset also shows the double occupancy of the 4-orbital HK model at $1/8$-hole-doping ($x=0.125$) and its comparison with QMC simulationsXu et al. (2024). Once again, the agreement is remarkable underscoring that orbital HK converges rapidly to Hubbard physics. Additionally, the computational cost for solving the 4-orbital HK model is considerably lower, both in terms of algorithm development (40 lines of Mathematica) and actual code execution, when compared to QMC or cluster-type DMFT simulations on the Hubbard model. Most crucially, the orbital HK model is more amenable to analytical calculations, an aspect we intend to explore in future studies.
The fact that the double occupancy only vanishes asymptotically as $U$ increases in the orbital HK model implies the presence of dynamical mixing between the upper and lower Mott sub-bands. This arises entirely from the non-commutativityHarris and Lange (1967); Meinders et al. (1993); Eskes et al. (1991) of the kinetic and potential energy and appears as $t/U$ corrections to the low-energy spectral weight (LESW). In the atomic limit, the LESW is strictly $2x$ (where $x$ is the doping level) because each hole can be occupied by a spin-up or spin-down electron. Any dynamical corrections to this necessarily generate double occupancy and hence increase the LESW strictly defined as
@@
In Fig. 4a, we compare the LESW of the 4-orbital HK model with Fig.(3) of Ref. [Meinders et al., 1993]. There is no qualitative difference with exact diagonalization on the (1D) Hubbard modelMeinders et al. (1993); Eskes et al. (1991). Both increase faster than $2x$. The semiconductor (dashed line) and Fermi liquid (dashed-dotted) results are shown for comparison. As the occupied part of the lower band has a weight $1-x$, the total weight in the lower band now exceeds $1+x$. As only $1+x$ electrons can occupy the lower band, dynamical spectral weight transfer (DSWT) defined as $\Lambda(x)-2x$ and plotted in Fig. (4b) impliesPhillips (2010) that the spectral weight in the lower band cannot be exhausted by counting electrons alone. Fig. (4b) is quantitatively in agreement with the Hubbard model (Fig. (4) of Ref. [Meinders et al., 1993]), the only difference being the maximum value which for 4-orbital HK is $0.2$ whereas it is $0.23$ for the (1D) Hubbard result. Consequently, that the LESW in the orbital HK model increases faster than $2x$ is a profoundly non-trivial result as no analytically solvable model has ever been formulated to capture this feature of the Hubbard model.
Figure 4: LESW and DSWT in the exact solution of the 4-orbital HK model ($U=10$) for the hopping parameters shown in panels (a) and (b) respectively. The solid line shown with slope $2x$ is the band HK or atomic Hubbard result. The dashed and dotted lines depict the semiconductor and Fermi liquid results respectively. Note, there is no qualitative difference with the exact diagonalization results for the Hubbard modelMeinders et al. (1993); Eskes et al. (1991). As $t/U$ increases, so does the DSWT.
All of these results point to a fundamental connection with the Hubbard model. Fig. 2 contains the essential idea. Because of the constraint, $N\times n=N_{\rm sites}$ and $\text{rBZ}=\text{BZ}/n$, $N$ the number of k-points, the rBZ is half the initial BZ when $n=2$, quartered for $n=4$, etc. as illustrated in Fig. 2a. Consequently, when $n=N_{\rm sites}$, no momentum degrees of freedom exist and only local orbitals are left, thereby leading to the $N_{\rm sites}$ Hubbard model. Alternatively, we can think of the orbital index as a label for the number of atoms per unit cell as depicted in (Fig. 2(b-d)). $N_{\rm sites}/n$, or $N$, is now the number of cells. Each cell contains the exact information of the $n$-site Hubbard model. For example, $n=2$ signifies that the orbital basis contains the information of the Hubbard dimer which has a non-degenerate ground state, and hence the spurious $2^{N}$ spin degeneracy of the band-HK model is removed. In addition to hybridization between orbitals in each cell, the cells are also connected in terms of hopping and interactions. For $n=4$, we show the explicit connection between the cells. The matrix $g({\bf k})$, representing the hopping connectivities in Fig. 2c, is given by
@@
for the 4-atom case, with $\varepsilon_{tx}=-2t\cos k_{x}$, $\varepsilon_{ty}=-2t\cos k_{y}$ and $\varepsilon_{t^{\prime}}=-4t^{\prime}\cos k_{x}\cos k_{y}$, identical in form to that in the $2\times 2$ Hubbard cluster with twisted boundary conditions (TBC) Huang (2020) where $\alpha$ is replaced with the site index $i$ and $k_{x}=\theta_{x}$ and $k_{y}=\theta_{y}$ where $\theta_{x}$ and $\theta_{y}$ define the TBC. Besides, TBC gives local-in-momentum interactions in the reduced BZ which is exactly the interaction term in Eq. (1). For any $n<N_{\rm sites}$, the orbital HK model corresponds to a $N_{\rm sites}/n$ copies of $n$-site Hubbard clusters all connected by TBC. The matrix that determines the TBC is purely local as it is determined by the dynamics. Hence, orbital HK is purely local in the rBZ regardless of the number of orbitals used, even $n=1$ corresponding to band HK whose locality is illustrated in Fig. 1. In this case, the TBC matrix is simply $\varepsilon_{\bf k}$. This exact mapping is a surprise and offers a new way of interpreting exact diagonalization with TBC. The most general statement is as follows. Theorem: The orbital HK model is equivalent to covering the Brillouin zone with $N_{\rm sites}/n$ Hubbard clusters each containing $n$-sites all connected via TBC. Clearly, the $N_{\rm sites}$-cluster Hubbard model results when $n=N_{\rm sites}$. Loosely, the $N_{\rm sites}/n$ identical copies throughout the Brillouin zone can be thought of as replica copies.
@@
Aside from DSWT, the orbital HK model also exhibits a pseudogap in the density of states. Shown in Fig. 5 is the density of states for the 4-orbital HK model with (Fig. 5(a,c)) and without (Fig. 5(b,d)) the next-nearest-neighbor hopping, $t^{\prime}$. Fig. 5(a) shows a suppression of the density of states at the chemical potential ($\rho(\omega=0)\lesssim 0.1$) in the underdoped region with $t^{\prime}=-0.25$ when $U\geq W$ ($W=8t$ is the bare bandwidth), indicative of a pseudogap in the absence of superconducting order. The complete tracking of the density of states at zero frequency in Fig. 5(c) displays the trend that as $U$ increases, the pseudo-gap region appears and extends to higher hole-doped density. In contrast, when $t^{\prime}=0$, the pseudo-gap suppression does not arise until $U\gg W$, as shown in Fig. 5(b,d). These observations are consistent with a recent studyHuang (2020) on a $2\times 2$ Hubbard cluster with TBC.
@@
Finally, we compute the heat capacity of the 4-orbital HK model at half-filling shown in Fig. 6. Most noticeable is the two-peak structure at $U>W$ representing a demarcation of the charge and spin degrees of freedom into high and low-temperature regimes, respectively, consistent with QMC simulations on the Hubbard model Duffy and Moreo (1997); Wang et al. (2022). Also at half-filling, we find a near-crossing of the heat capacity curves as a function of $U$ at a temperature intermediate between the spin and charge excitations. The Maxwell relations governing the entropy dictateVollhardt (1997) such a crossing. Since there is no sign-problem restriction, we can access the low-temperature heat capacity exactly. As we show in Fig. 6(b), the low-temperature heat capacity data follow a power-law increase detailed in the supplement. While such algebraic growth might seem counterintuitive for a Mott insulator, the first-excited state is charge neutralMai et al. (2024); Zhao et al. (2023a) as shown in the supplement. Such charge-neutral excitations determine the low-T behavior of the specific heat.